# Neural Geometry Processing via Spherical Neural Surfaces

### v2

@@ Abstract
Neural surfaces (e.g., neural map encoding, deep implicits and neural radiance fields) have recently gained popularity because of their generic structure (e.g., multi-layer perceptron) and easy integration with modern learning-based setups. Traditionally, we have a rich toolbox of geometry processing algorithms designed for polygonal meshes to analyze and operate on surface geometry. In the absence of an analogous toolbox, neural representations are typically discretized and converted into a mesh, before applying any geometry processing algorithm. This is unsatisfactory and, as we demonstrate, unnecessary. In this work, we propose a spherical neural surface representation for genus-0 surfaces and demonstrate how to compute core geometric operators directly on this representation. Namely, we estimate surface normals and first and second fundamental forms of the surface, as well as compute surface gradient, surface divergence and Laplace Beltrami operator on scalar/vector fields defined on the surface. Our representation is fully seamless, overcoming a key limitation of similar explicit representations such as Neural Surface Maps [MAKM21]. These operators, in turn, enable geometry processing directly on the neural representations without any unnecessary meshing. We demonstrate illustrative applications in (neural) spectral analysis, heat flow and mean curvature flow, and evaluate robustness to isometric shape variations. We propose theoretical formulations and validate their numerical estimates, against analytical estimates, mesh-based baselines, and neural alternatives, where available. By systematically linking neural surface representations with classical geometry processing algorithms, we believe that this work can become a key ingredient in enabling neural geometry processing. Code will be released upon acceptance, accessible from the project webpage.
@@ 1 Introduction
We propose a different approach to discretization-free geometry processing. First, we create a smooth and differentiable representation — spherical neural surfaces (SNS) — where we encode individual genus-0 surfaces using dedicated neural networks. Having a differentiable encoding allows one to directly compute surface normals, tangent planes, and Fundamental Forms directly with continuous differential geometry. The First Fundamental Form allows us to compute attributes of the parameterization, such as local area distortion. The Second Fundamental form allows us to compute surface attributes such as mean curvature, Gauss curvature, and principal curvature directions. Essentially, encoding surfaces as seamless neural maps unlocks definitions from differential geometry.
@@
    compute surface normals, First and Second Fundamental Forms, curvature, continuous Laplace Beltrami operators, and their smallest spectral modes without using any unnecessary discretization;
@@ 2 Related Works
Estimating differential quantities.
@@
Laplace Beltrami operators on meshes.
The Laplace-Beltrami operator (LBO) is an indispensable tool in spectral geometry processing, as it is used to analyze and manipulate the intrinsic properties of shapes. Spectral mesh processing [LZ10] leverages the eigenvalues and eigenfunctions of the LBO for tasks such as mesh smoothing and shape segmentation. Many subsequent efforts have demonstrated the use of LBO towards shape analysis and understanding (e.g., shapeDNA [RWP06], global point signatures [Rus07], heat kernel signature [SOG09]), eventually culminating into the functional map framework [OBCS12]. An earlier signal processing framework [Tau95] has been revisited with the LBO operator [LZ10] for multi-scale processing applications (e.g., smoothing and feature enhancement). Although the Laplace-Beltrami operator on triangular meshes or, more generally, on polyhedral meshes, can be computed, these operators depend on the underlying discretization of the surface (i.e., placement of vertices and their connectivity). See [BAB23] for a discussion. In the case of triangular meshes, the commonly used discretizations are the Uniform LBO and the cotan LBO. In Section 6, we compare our neural LBO with the cotan discretization and discuss their relative merits.
@@ 3.1 Overfitting an SNS to a Triangle Mesh
$\mathcal{L}_{\textit{normal}}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{n}_{S_{\theta}}({p_{i}})-\mathbf{n}_{\textit{mesh}}({p_{i}})\right\|^{2}=\frac{2}{N}\sum_{i=1}^{N}(1-\cos{\alpha_{i}(p_{i})}).$ (2)
In this expression, $\mathbf{n}_{\textit{mesh}}({p_{i}})$ is the (outwards) unit-normal on the mesh, and $\mathbf{n}_{S_{\theta}}({p_{i}})$ is the corresponding (outwards) unit-normal of the smooth parametrization at $p_{i}$, which is derived analytically from the Jacobian of $S_{\theta}$ at $p_{i}$ (see Section 3.2); the angle $\alpha(p_{i})$ is the angle between $\mathbf{n}_{\textit{mesh}}({p_{i}})$ and $\mathbf{n}_{S_{\theta}}({p_{i}})$. Figure 1 shows an example of an SNS over training stages.
@@ 3.2 Computing Differential Quantities using SNS
and then turn this into a $3\times 2$ Jacobian in local co-ordinates, by composing it with a $3\times 2$ orthonormal matrix ($R$):
@@
The matrix $R$ maps vectors in 2D local co-ordinates to tangent vectors on the sphere at a particular point, and the matrix $J(S)$ maps tangent vectors on the sphere to tangent vectors on the surface $S$. If $u$ and $v$ are the corresponding coordinates on the tangent plane of the sphere, then the columns of $\mathbf{J}^{\textit{local}}_{S}$ are equal to the partial derivatives $S_{u}$ and $S_{v}$ respectively:
@@
This matrix is an orthonormal version of the usual spherical polar coordinates Jacobian. We avoid degeneracy at the poles by selectively re-labeling the $x$, $y$, and $z$ coordinates so that $\theta$ is never close to zero or $\pi$. (This is equivalent to choosing a different chart for the points within some radius of the poles.)
@@
To compute the Second Fundamental Form, we first compute the second partial derivatives of the parametrization with respect to the local co-ordinates, and then find the dot products of the second derivatives with the normal, $\mathbf{n}=\mathbf{n}_{S_{\theta}}$. The second derivatives are:
@@
Figure 2: First and Second Fundamental Forms. We show an analytically defined surface, spike (bottom) and an SNS (top row) that has been overfitted to a mesh of the surface (with 40962 vertices). We show various differential quantities computed symbolically using Matlab (bottom) and computed on the SNS using our method (top). Our results align very closely with the analytically-computed results.
@@ 4.1 Laplace-Beltrami Operator on Smooth Surfaces
$\Delta f:=\nabla\cdot\nabla f=\sum_{i=1}^{n}\frac{\partial^{2}f}{\partial^{2}x_{i}}.$ (12)
@@
where $\mathbf{n}$ is the unit surface normal at the point. By choosing a suitable co-ordinate frame, one can show that this definition is independent of the choice of extension. The surface divergence of the vector field $\mathbf{F}$ is defined as
@@
As in section 3.2, these formulae are easily implemented for functions defined on an SNS. All that is required to implement equation 16 is that we have access to an extension $\tilde{f}$ and its first and second derivatives. Therefore it can be applied when ${f}$ is (the restriction to the surface of) a closed-form function on the embedding space (e.g. figures 10 and 11), or when $f$ and $\tilde{f}$ are defined by a network, allowing us to use autodiff (see section 4.2).
We have provided Equation 16) as an alternative formula to the standard definition (Equation 13) because where a field evolves on a domain that remains constant, it allows us to precompute the $\mathbf{H}$ and $\mathbf{n}$ and then only the Euclidean quantities ($\nabla\tilde{f}$, $\mathbf{H}(\tilde{f})$, and $\Delta\tilde{f}$) need to be updated. In Figure 8 (2nd and 3rd row) we demonstrate both methods and there is no visible difference.
@@ 4.2 Spectrum of the Laplace-Beltrami Operator
Given two scalar functions, $f$ and $g$, defined on the same regular surface $\Sigma$, we can define their inner product to be,
@@
Then, the eigenfunctions of $\Delta_{\Sigma}$ are the minimizers of the Rayleigh Quotient, as given by,
@@
We use $\Psi_{0}$ to represent the first eigenfunction, which is always a constant function. The constraint states that the eigenfunctions are orthogonal inside the inner-product space $L^{2}(\Sigma)$.
In addition, the Rayleigh Quotient of the eigenfunction $\Psi_{i}$ is the positive of the corresponding eigenvalue, i.e.,
@@
Figure 4: Encoding Scalar Fields on Neural Surfaces. We encode scalar fields on the sphere (and corresponding extensions into $\mathbb{R}^{3}$) as MLPs $g_{\eta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, with parameters $\eta$. Because $S_{\theta}$, considered as a parametrization, is bijective, then we can implicitly define any smooth scalar field $h$ on the surface in terms of one of the functions $g_{\eta}$ (restricted to $\mathbb{S}^{2}$). We visualize the one-to-one correspondence between scalar fields defined on the surface, and scalar fields defined on the sphere, by using vertex-colors to display the fields.
@@
Figure 6: Accuracy of Differential Quantity Estimation. As shown by the colourmaps, the quality of the estimation of differential quantities (mean curvature, Gauss curvature, minimum curvature direction and normals) decreases gradually as mesh-resolution decreases. We show results for an SNS overfitted to analytical parametrisation, with 102400 sample points (top), and SNSs fitted to meshes with 40962, 10242 and 2562 vertices. (The errors are sampled at a much higher resolution, at 655362 points.) The level of error for the finest level mesh is close to the level of error for the ‘analytical’ SNS and we argue that this demonstrates the ‘interpolation’ capability of our method: it shows that fitting an SNS to an appropriately high-resolution mesh successfully interpolates the underlying surface (even for meshes where the analytic parametrisation is not available). We also show plots of the error on the various differential quantities, compared against three additional methods (when functions are available): Monge-fitting, NFGP and i3d. A minimum curvature direction function was not available for NFGP.
@@ 5 Spectral Analysis using SNS
To find the continuous eigenmodes of the Laplace-Beltrami operator on an SNS, we require a continuous and differentiable representation for scalar functions defined on the surface. In our framework, we represent smooth scalar fields on the sphere $\mathbb{S}^{2}$ by MLPs, $g_{\eta}$, from $\mathbb{R}^{3}$ to $\mathbb{R}$, which are parametrized by $\eta$. We only ever evaluate the MLP $g_{\eta}$ on $\mathbb{S}^{2}\in\mathbb{R}^{3}$, but because the network’s domain is $\mathbb{R}^{3}$ then it defines an extension of the scalar field, and this allows us to compute the Euclidean gradient $\nabla g_{\eta}$. Then, if $S_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is an SNS, any smooth scalar field $h$ can be defined implicitly by the equation
@@
where we sequentially compute the eigenfunctions, and $h_{i}$ is the $i$th smallest eigenfunction. Recall that we represent each such eigenfunction using a dedicated MLP (except for $h_{0}$, which is the constant eigenfunction). The loss term $\mathcal{L}_{ortho}$ encourages the scalar field $h$ to be orthogonal to all previously-found eigenfunctions (including the constant eigenfunction). The quantity $Q_{S_{\theta}}(h)$ depends on the Jacobian of $S_{\theta}$, through $\nabla_{S_{\theta}}h$ and $\nabla h$. However, since the surface does not change during optimization and $\theta$ stays fixed, we precompute $(\mathbf{J}(S_{\theta})^{T})^{-1}$. Note that, if we were to optimize over the subspace of the unit-norm functions in $L^{2}(\Sigma)$, then we could replace the Rayleigh Quotient with the Dirichlet energy. However, it was a more natural design choice for us to parametrize the entire space $L^{2}(\Sigma)$ and optimize the Rayleigh Quotient, rather than strictly constrain the functions to have unit-norm and minimize the Dirichlet Energy.
@@
Finally, since we cannot exactly compute the inner products in these loss terms, we approximate the inner products by Monte-Carlo sampling. Specifically, we approximate,
$f,g\right>_{L^{2}(S)}\approx\frac{Area(S)}{N}\sum_{i=1}^{N}f(p_{i})g(p_{i}),$ (28)
where $\{p_{i}\}_{1\leq i\leq N}$ is a set of uniformly distributed points on the surface $S_{\theta}$; $f(p_{i})$ and $g(p_{i})$ denote calls to the scalar field MLPs. (In the overfitting stage, we normalize all our Spherical Neural Surfaces to have an area equal to $4\pi$.) For the Monte Carlo step, we generate $N$ uniformly distributed points on the surface $S_{\theta}$ via the following steps (rejection sampling):
@@
    Generate $M\gg N_{target}$ samples ($p_{i}$) from the 3D normal distribution $\mathcal{N}(\mathbf{\mu}=\mathbf{0},{\Sigma}=\mathbf{I})$. Normalize, to unit norm, to produce a dense uniform distribution of points on the sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$.
@@
    Compute the local area distortion ($d_{i}=\sqrt{E_{i}G_{i}-F_{i}^{2}}$), for each point $p_{i}$, using the First Fundamental Form.
@@
    Perform rejection sampling, such that the probability of keeping point $p_{j}$ is equal to $\frac{d_{j}}{\sum_{i=1}^{M}d_{i}}N_{target}$.

### v3

@@ Abstract
Neural surfaces (e.g., neural map encoding, deep implicit, and neural radiance fields) have recently gained popularity because of their generic structure (e.g., multi-layer perceptron) and easy integration with modern learning-based setups. Traditionally, we have a rich toolbox of geometry processing algorithms designed for polygonal meshes to analyze and operate on surface geometry. Without an analogous toolbox, neural representations are typically discretized and converted into a mesh, before applying any geometry processing algorithm. This is unsatisfactory and, as we demonstrate, unnecessary. In this work, we propose a spherical neural surface representation for genus-0 surfaces and demonstrate how to compute core geometric operators directly on this representation. Namely, we estimate surface normals and first and second fundamental forms of the surface, as well as compute surface gradient, surface divergence and Laplace Beltrami operator on scalar/vector fields defined on the surface. Our representation is fully seamless, overcoming a key limitation of similar explicit representations such as Neural Surface Maps [MAKM21]. These operators, in turn, enable geometry processing directly on the neural representations without any unnecessary meshing. We demonstrate illustrative applications in (neural) spectral analysis, heat flow and mean curvature flow, and evaluate robustness to isometric shape variations. We propose theoretical formulations and validate their numerical estimates, against analytical estimates, mesh-based baselines, and neural alternatives, where available. By systematically linking neural surface representations with classical geometry processing algorithms, we believe this work can become a key ingredient in enabling neural geometry processing. Code is available via the project webpage.
volume: 44issue: 2
@@ 1 Introduction
We propose a different approach to discretization-free geometry processing. First, we create a smooth and differentiable representation — spherical neural surfaces (SNS) — where we encode individual genus-0 surfaces using dedicated neural networks. Having a differentiable encoding allows one to directly compute surface normals, tangent planes, and Fundamental Forms directly with continuous differential geometry. The First Fundamental Form allows us to compute attributes of the parametrization, such as local area distortion. The Second Fundamental form allows us to compute surface attributes such as mean curvature, Gauss curvature, and principal curvature directions. Essentially, encoding surfaces as seamless neural maps unlocks definitions from differential geometry.
@@
    compute surface normals, First and Second Fundamental Forms, curvature, continuous Laplace Beltrami operators without unnecessary discretization;
@@
    approximate the smallest spectral modes of the Laplace Beltrami operator on spherical neural surfaces;
  4. (iv)
@@ 2 Related Works
Estimating Differential Quantities
@@
Laplace Beltrami Operators on Meshes
The Laplace-Beltrami operator (LBO) is an indispensable tool in spectral geometry processing, as it is used to analyze and manipulate the intrinsic properties of shapes. Spectral mesh processing [LZ10] leverages the eigenvalues and eigenfunctions of the LBO for tasks such as mesh smoothing and shape segmentation. Many subsequent efforts have demonstrated the use of LBO towards shape analysis and understanding (e.g., shapeDNA [RWP06], global point signatures [Rus07], heat kernel signature [SOG09]), eventually culminating into the functional map framework [OBCS12]. An earlier signal processing framework [Tau95] has been revisited with the LBO operator [LZ10] for multi-scale processing applications (e.g., smoothing and feature enhancement). Although the Laplace-Beltrami operator on triangular meshes or, more generally, on polyhedral meshes, can be computed, these operators depend on the underlying discretization of the surface (i.e., placement of vertices and their connectivity). See [BAB23] for a discussion. In the case of triangular meshes, the commonly used discretizations are the Uniform Laplace Operator and the cotan LBO. In Section 6, we compare our neural LBO with the cotan discretization and discuss their relative merits.
@@ 3.1 Overfitting an SNS to a Triangle Mesh
$\mathcal{L}_{\textit{normal}}:=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{n}_{S_{\theta}}({p_{i}})-\mathbf{n}_{\textit{mesh}}({q_{i}})\right\|^{2}=\frac{2}{N}\sum_{i=1}^{N}(1-\cos{\alpha_{i}}).$ (2)
In this expression, $\mathbf{n}_{\textit{mesh}}({q_{i}})$ is the (outwards) unit-normal on the mesh at $q_{i}$, and $\mathbf{n}_{S_{\theta}}({p_{i}})$ is the corresponding (outwards) unit-normal of the smooth parametrization at $S_{\theta}(p_{i})$, which is derived analytically from the Jacobian of $S_{\theta}$ at $p_{i}$ (see Section 3.2); the angle $\alpha_{i}$ is the angle between $\mathbf{n}_{\textit{mesh}}({q_{i}})$ and $\mathbf{n}_{S_{\theta}}({p_{i}})$. Figure 1 shows an example of an SNS over training stages.
@@ 3.2 Computing Differential Quantities using SNS
and then turn this into a $3\times 2$ Jacobian in local coordinates, by composing it with a $3\times 2$ orthonormal matrix ($R$):
@@
The matrix $R$ maps vectors in 2D local coordinates to tangent vectors on the sphere at a particular point, and the matrix $\mathbf{J}(S)$ maps tangent vectors on the sphere to tangent vectors on the surface $S$. If $u$ and $v$ are the local coordinates on the tangent plane of the sphere, then the columns of $\mathbf{J}^{\textit{local}}_{S}$ are equal to the partial derivatives $S_{u}$ and $S_{v}$ respectively:
@@
The matrix $R$ is an orthonormal version of the usual spherical polar coordinates Jacobian. We avoid degeneracy at the poles by selectively re-labeling the $x$, $y$, and $z$ coordinates so that $\theta$ is never close to zero or $\pi$. (This is equivalent to choosing a different chart for the points within some radius of the poles.)
@@
To compute the Second Fundamental Form, we first compute the second partial derivatives of the parametrization with respect to the local coordinates, and then find the dot products of the second derivatives with the normal, $\mathbf{n}=\mathbf{n}_{S_{\theta}}$. The second derivatives are:
@@
Figure 2: First and Second Fundamental Forms. We show an analytically defined surface, star (bottom) and an SNS (top row) that has been overfitted to a mesh of the surface (with 40962 vertices). We show various differential quantities computed symbolically using Matlab (bottom) and computed on the SNS using our method (top). Our results align very closely with the analytically-computed results.
@@ 4.1 Laplace-Beltrami Operator on Smooth Surfaces
$\Delta f:=\nabla\cdot\nabla f=\sum_{i=1}^{n}\frac{\partial^{2}f}{\partial x_{i}^{2}}.$ (12)
@@
where $\mathbf{n}$ is the unit surface normal at the point. One can show that this definition is independent of the choice of extension. The surface divergence of the vector field $\mathbf{F}$ is defined as
@@
As in Section 3.2, these formulae are easily implemented for functions defined on an SNS. All that is required to implement Equation 16 is that we have access to an extension $\tilde{f}$ and its first and second derivatives. Therefore it can be applied when ${f}$ is (the restriction to the surface of) a closed-form function on the embedding space (e.g., Figures 10 and 11), or when $f$ and $\tilde{f}$ are defined by a network, allowing us to use autodiff (see Section 4.2).
We have provided Equation 16 as an alternative formula to the standard definition (Equation 13) because where a field evolves on a domain that remains constant, it allows us to precompute the mean curvature $H$ and normal $\mathbf{n}$, and then only the Euclidean quantities ($\nabla\tilde{f}$, $\mathbf{H}(\tilde{f})$, and $\Delta\tilde{f}$) need to be updated. In Figure 8 (2nd and 3rd row) we demonstrate both methods and there is no visible difference.
@@ 4.2 Spectrum of the Laplace-Beltrami Operator
Given two scalar functions, $f$ and $g$, defined on the same regular surface $\Sigma$, their $L^{2}$ inner product is
@@
Then, the first $k$ eigenfunctions of $\Delta_{\Sigma}$ are the minimizers of the Rayleigh Quotient, as given by,
@@
We use $\Psi_{0}$ to denote the first eigenfunction, which is always a constant function. The constraint states that the eigenfunctions are orthogonal inside the inner-product space $L^{2}(\Sigma)$.
In addition, the Rayleigh Quotient of the eigenfunction $\Psi_{i}$ is the positive of the corresponding (negative) eigenvalue, i.e.,
@@
Figure 4: Encoding Scalar Fields on Neural Surfaces. We encode scalar fields on the sphere (and corresponding extensions into $\mathbb{R}^{3}$) as MLPs $g_{\eta}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, with parameters $\eta$. Because $S_{\theta}$, considered as a parametrization, is bijective, then we can implicitly define any smooth scalar field $h$ on the surface in terms of one of the functions $g_{\eta}$ (restricted to $\mathbb{S}^{2}$). We visualize the one-to-one correspondence between scalar fields defined on the surface, and scalar fields defined on the sphere, by using vertex-colours to display the fields.
@@
Figure 6: Accuracy of Differential Quantity Estimation. As shown by the colourmaps, the quality of the estimation of differential quantities (mean curvature, Gauss curvature, minimum curvature direction and normals) decreases gradually as mesh-resolution decreases. We show results for an SNS overfitted to analytical parametrization, with 102400 sample points (top), and SNSs fitted to meshes with 40962, 10242 and 2562 vertices. (The errors are sampled at a much higher resolution, at 655362 points.) The level of error for the finest level mesh is close to the level of error for the ‘analytical’ SNS and we argue that this demonstrates the ‘interpolation’ capability of our method: it shows that fitting an SNS to an appropriately high-resolution mesh successfully interpolates the underlying surface (even for meshes where the analytic parametrization is not available). We also show plots of the error on the various differential quantities, compared against three additional methods (when functions are available): Monge-fitting, NFGP and i3d. A minimum curvature direction function was not available for NFGP.
@@ 5 Spectral Analysis using SNS
To find the continuous eigenmodes of the Laplace-Beltrami operator on an SNS, we require a continuous and differentiable representation for scalar functions defined on the surface. In our framework, we represent smooth scalar fields on the sphere $\mathbb{S}^{2}$ by MLPs, $g_{\eta}$, from $\mathbb{R}^{3}$ to $\mathbb{R}$, which are parametrized by $\eta$. We only ever evaluate the MLP $g_{\eta}$ on $\mathbb{S}^{2}\in\mathbb{R}^{3}$, but because the network’s domain is $\mathbb{R}^{3}$ then it defines an extension of the scalar field, and this allows us to compute the Euclidean gradient $\nabla g_{\eta}$. Then, if $S_{\theta}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is an SNS, any smooth scalar field $h$ whose domain is the surface defined by $S_{\Theta}$ can be defined implicitly by the equation
@@
where we sequentially compute the eigenfunctions, and $h_{i}$ is the $i$th smallest eigenfunction. Recall that we represent each such eigenfunction using a dedicated MLP (except for $h_{0}$, which is the constant eigenfunction). The loss term $\mathcal{L}_{ortho}$ encourages the scalar field $h$ to be orthogonal to all previously-found eigenfunctions (including the constant eigenfunction). The quantity $Q_{S_{\theta}}(h)$ depends on the Jacobian of $S_{\theta}$, through $\nabla_{S_{\theta}}h$ and $\nabla h$. However, since the surface does not change during optimization and $\theta$ stays fixed, we precompute $(\mathbf{J}(S_{\theta})^{T})^{-1}$. Note that, if we were to optimize over the subspace of the unit-norm functions in $L^{2}(\Sigma)$, then we could replace the Rayleigh Quotient with the Dirichlet energy. However, it is difficult to construct a function, based on an MLP, that produces only unit-norm functions (in the $L^{2}$ sense). In initial experiments, we used a unit-norm loss in combination with the Dirichlet Energy and orthogonality loss, but we observed some robustness issues, such as a tendency for the functions to collapse to zero. We found that robustness increased when we replaced the Dirichlet Energy by the Rayleigh Quotient, and we think that this is the Rayleigh Quotient, by cancelling out the effect of magnitude, is less in opposition to the unit-norm loss than the Dirichlet energy is. The Rayleigh Quotient design choice greatly reduces the importance of the unit-norm loss, but we retained it as a regularizer, to prevent functions from collapsing to zero.
@@
Finally, since we cannot exactly compute the integrals for the inner products in these loss terms, we approximate the inner products using Monte-Carlo sampling. Specifically, we approximate,
$h,h_{i}\right>_{L^{2}(S)}\approx\frac{Area(S)}{N}\sum_{j=1}^{N}h(q_{j})h_{i}(q_{j}),$ (28)
where $\{q_{j}\}_{1\leq j\leq N}$ is a set of uniformly distributed points on the surface $S_{\theta}$, corresponding to points $\{p_{j}\}_{1\leq j\leq N}$ on the sphere; $h(q_{j})$ and $h_{i}(q_{j})$ are evaluated by applying scalar field MLPs $g_{\eta}$ and $g_{\eta_{i}}$ to $p_{j}$. (In the overfitting stage, we normalize all our Spherical Neural Surfaces to have an area equal to $4\pi$.) For the Monte Carlo step, we generate $N$ uniformly distributed points on the surface $S_{\theta}$ via the following steps (rejection sampling):
@@
    Generate $M\gg N_{target}$ samples ($p_{j}$) from the 3D normal distribution $\mathcal{N}(\mathbf{\mu}=\mathbf{0},{\Sigma}=\mathbf{I})$. Normalize, to unit norm, to produce a dense uniform distribution of points on the sphere $\mathbb{S}^{2}\subset\mathbb{R}^{3}$.
@@
    Compute the local area distortion ($d_{j}=\sqrt{E_{j}G_{j}-F_{j}^{2}}$), for each point $p_{j}$, using the First Fundamental Form.
@@
    Perform rejection sampling, such that the probability of keeping point $p_{k}$ is equal to $\frac{d_{k}}{\sum_{j=1}^{M}d_{j}}N_{target}$.
  4. (iv)
    Push the samples $p_{j}$ (on the sphere) through the map $S_{\theta}$.